in young's two-slit experiment using light of wavelength 6.0×10^-7m, the slits were 0.40mm apart and the distance of the slits to the screen was 1.20m find the separation of the fringes. what is the angle in radians subtended by a central pair of bright fringes at the slit? 

1
Expert's answer
2021-12-14T15:09:05-0500
"d\sin\theta=m\lambda, \\\space\\
\theta=\arcsin\frac{m\lambda}{d}=1.5·10^{-3},\\\space\\
\phi=2\theta=0.003\text{ rad}."